# D-NPC: Dynamic Neural Point Clouds for Non-Rigid View Synthesis from Monocular Video

### v1

@@ 4.2 Qualitative Results
TiNeuVox RoDynRF 4D-GS Ours Ground Truth
[…]
Figure 3: Comparison on the iPhone dataset. We compare on two challenging scenes of the iPhone dataset. Our method better preserves the object shapes and fine details despite complex deformations. Methods indicated by $\dagger$ are the highest quality versions provided by Gao et al. [15]
We show qualitative results on the NVIDIA (Fig. 2) and iPhone (Fig. 3) datasets. For comparisons on the iPhone sequences, we show the best version of HyperNeRF [35] and T-NeRF [37] according to image quality metrics, using all additional regularizers (Lidar depth, random background, and distortion loss) proposed by Gao et al. [15]. For our method, we report values for two versions, trained with the provided Lidar or monocular depth data. Again, we find that our method achieves a competitive and oftentimes superior perceived image quality, preserving higher frequencies and clean edges. For dynamic view synthesis results, please see our supplemental video.
@@ 4.3 Ablation Study
Table 3: Comparison of ablated D-NPC versions in terms of image quality, model size and speed.
@@ 5 Discussion and Limitations
In our experiments, our D-NPC yields the best trade-off between optimization time and image quality and is currently the only monocular method providing real-time frame rates for interactive applications. Especially in terms of perceptual LPIPS and high image frequencies, our method quantitatively and visually outperforms recent monocular view synthesis methods while only requiring a fraction of the optimization time. However, there is room for improvement: Our method relies on a decent quality of camera poses and monocular depth. Moreover, our initialization process assumes that the dynamic content is fully contained in the camera frustum, resulting in visible crops when synthesizing far-off novel views, as shown in Fig. 4.
@@ 6 Conclusion
We presented Dynamic Neural Point Clouds, an efficient implicit point cloud-driven [17] approach for non-rigid novel-view synthesis from monocular video inputs. In contrast to existing ray marching-based NeRF methods, our rasterization-based forward rendering approach enables a direct and elegant way to initialize and guide the reconstruction from data-driven priors, not only making it one of the fastest approaches in terms of optimization time, but also enabling rendering at real-time frame rates. While our experiments show our novel approach already yields competitive image quality according to several metrics on common benchmark datasets, we believe that our work constitutes a promising new direction for future research and extensions, like the integration of diffusion-based priors.
Acknowledgments and Disclosure of Funding
We thank Jann-Ole Henningson for help with figures and writing.
This work was partially funded by the German Research Foundation (DFG) projects “Immersive Digital Reality” (ID 283369270) and “Real-Action VR” (ID 523421583).
Appendix A Appendix – Supplemental Material
In this supplemental, we provide details on our model configuration (Sec. A.1) and optimization procedure (Sec. A.2), include additional results (Sec. A.3), and discuss concurrent work on arXiv (Sec. A.4) as well as the societal impact (Sec. A.5) of our work.
A.1 Model Details
Table 4: Appearance Field Hash Grid Configurations.
Parameter 3D Static Grid 4D Dynamic Grid
Base Resolution 16 16
Resolution Levels 10 8
Per Level Scale 2 2
Features Per Level 4 4
Logarithmic Size 21 22
In our model, the crucial neural components regarding the image quality are the appearance feature grids and the neural rendering network. To model our static and dynamic appearance fields, we use the fast multi-resolution hash encoding implementation by Müller et al. [33, 32] After initially carving our sampling distribution voxel grid using the monocular depth and dynamic segmentation priors, we set the spatial scale of our 3D and 4D feature grids to the minimum axis-aligned bounding box enclosing all remaining static and dynamic voxels, respectively. We parameterize the hash-grids according to the values in Tab. 4, giving more total capacity to the higher dimensional dynamic representation. Both hash encodings are followed by a shallow ReLU-activated MLP with one hidden $64$-neuron layer to resolve hash collisions.
For our neural rendering network used for hole-filling, we apply a simple image-to-image translation approach based on a U-Net architecture [38]. In contrast to Hahlbohm et al. [17], we do not use any multi-resolution inputs. Instead, we use two downsampling and consecutive upsampling blocks, each consisting of two ReLU-activated 2D convolutions with kernel size $3$, followed by average pooling / bilinear upsampling. During down- and upscaling, each convolution block doubles/halves the amount of feature channels while adjusting the spatial resolution, going from an initial $64$ channels at the finest to a maximum of $256$ channels at the coarsest spatial resolution.
A.2 Optimization Details
Table 5: Loss Hyperparameters.
Hyperparameter Value
$\lambda_{\text{photo, Cauchy}}$ 1.0
$\lambda_{\text{photo, D-SSIM}}$ 0.9
$\lambda_{\text{photo, LPIPS}}$ 0.05
$\lambda_{\text{depth}}$ 20.0
$\lambda_{\text{seg}}$ $1\mathrm{e}{-3}$
$\lambda_{\text{dist}}$ 1.0
$k$ 0.5
We train our model for a total of 10K iterations on a consumer-grade RTX 4090 GPU. During each iteration, we sample an explicit point cloud consisting of $4$M points (i.e., a batch of spatial 3D locations). While the sampling distribution is manually updated based on statistics of the extracted explicit point clouds (see Sec. 3.2.1), the differentiable rasterizer and neural rendering network enable end-to-end optimization using gradient descent for the network and hash-grid parameters. We use the Adam [20] optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.99$, $\epsilon=1\mathrm{e}{-15}$) with an initial learning rate of $1\mathrm{e}{-2}$ for the appearance hash-grids and $5\mathrm{e}{-4}$ for the rendering network, which are exponentially decayed to $\frac{1}{3}\mathrm{e}{-3}$ and $5\mathrm{e}{-5}$ during optimization. For our loss hyperparameters (Sec. 3.4), we use the values provided in Tab. 5. Note that, after initially training with a high value $\lambda_{depth}$ to guide the optimization in a reasonable direction, we turn off our depth loss after $500$ training iterations to allow the model to recover from errors in the monocular depth estimates. In contrast to D2NeRF, who use $k=2$ to bias the binary segmentation loss towards static background, which usually constitutes a majority of the scene, we use a value of $0.5$ to favor the dynamic model. This is due to the fact that our initialization procedure already discards large amounts of dynamic model evaluation in static areas. As a result, the remaining dynamic samples are already biased towards potentially dynamic areas, which we reflect in the choice of skew $k$. In addition to the losses already discussed in Sec. 3.4, we further apply weight decay (i.e., an L2 loss) with a factor of $2.5$ on the appearance grid parameters [5]. In practice, we find that this decay helps to prevent overfitting and improves the quality of generated novel views. Our Python / CUDA implementation will be made publicly available.
A.3 Additional Results
We provide detailed breakdowns on the NVIDIA and iPhone datasets in Tab. 6 and Tab. 7, respectively. As evident from the per-scene results, our method consistently achieves competitive results across a variety of challenging scenes, and particularly excels at perceptual LPIPS scores due to its ability to retain high-frequency details.
Table 6: Quantitative comparison on the NVIDIA dataset. We compare our method against recent dynamic view synthesis approaches on the NVIDIA dynamic scenes dataset. Our method achieves the best scores in terms of perceptual image quality. The three best results are highlighted in green in descending order of saturation.
Balloon1 Balloon2 Jumping Playground
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
T-NeRF [37] 18.54 0.448 0.416 20.69 0.573 0.335 18.04 0.496 0.541 14.68 0.240 0.527
D-NeRF [37] 19.06 0.492 0.382 20.76 0.557 0.367 22.36 0.712 0.318 20.18 0.670 0.250
HyperNeRF [35] 13.96 0.214 0.677 16.57 0.286 0.544 18.34 0.521 0.460 13.17 0.179 0.603
NR-NeRF [49] 17.39 0.369 0.497 22.41 0.703 0.314 20.09 0.644 0.418 15.06 0.248 0.450
NSFF [25] 21.96 0.701 0.281 24.27 0.731 0.276 24.65 0.813 0.227 21.22 0.705 0.263
DynamicNeRF [14] 22.36 0.775 0.169 27.06 0.859 0.113 24.68 0.842 0.153 24.15 0.849 0.145
RoDynRF [27] 22.37 0.782 0.157 26.19 0.846 0.107 25.66 0.853 0.129 24.96 0.899 0.080
TiNeuVox [11] 17.30 0.392 0.500 19.06 0.458 0.396 20.81 0.641 0.399 13.84 0.213 0.541
4D-GS [55] 20.82 0.666 0.271 24.34 0.775 0.211 22.16 0.735 0.270 20.22 0.724 0.192
Ours 22.26 0.779 0.161 25.84 0.835 0.107 24.51 0.839 0.116 23.59 0.866 0.099
Skating Truck Umbrella Average
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
T-NeRF [37] 20.32 0.551 0.562 18.33 0.462 0.523 17.69 0.284 0.670 18.33 0.436 0.511
D-NeRF [37] 22.48 0.678 0.403 24.10 0.743 0.213 21.47 0.554 0.310 21.49 0.629 0.320
HyperNeRF [35] 21.97 0.640 0.322 20.61 0.493 0.394 18.59 0.308 0.517 17.60 0.377 0.502
NR-NeRF [49] 23.95 0.745 0.342 19.33 0.485 0.548 19.63 0.376 0.471 19.69 0.510 0.434
NSFF [25] 29.29 0.889 0.186 25.96 0.775 0.248 22.97 0.644 0.315 24.33 0.751 0.257
DynamicNeRF [14] 32.66 0.951 0.077 28.56 0.872 0.146 23.26 0.711 0.190 26.10 0.837 0.142
RoDynRF [27] 28.68 0.939 0.076 29.13 0.900 0.089 24.26 0.757 0.135 25.89 0.854 0.110
TiNeuVox [11] 23.32 0.723 0.283 23.86 0.662 0.328 20.00 0.416 0.439 19.74 0.501 0.412
4D-GS [55] 26.04 0.876 0.171 25.08 0.753 0.233 21.59 0.588 0.309 22.89 0.731 0.237
Ours 30.22 0.948 0.061 28.92 0.897 0.084 24.15 0.747 0.136 25.64 0.845 0.109
Table 7: Quantitative evaluation on the iPhone dataset. We compare our method against related approaches using masked metrics based on co-visibility [15]. Methods marked with $\dagger$ are trained using the code and additional regularizers introduced by Gao et al. [15]. We further show values for our method trained on monocular depth instead of the provided Lidar data. The three best results are highlighted in green in descending order of saturation.
Apple Block Paper-windmill Space-out
Method mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$
HyperNeRF+B+D+S [35] 17.64 0.743 0.478 17.54 0.670 0.331 17.38 0.382 0.209 17.93 0.605 0.320
Nerfies+B+D+S [34] 17.54 0.750 0.478 16.61 0.639 0.389 17.34 0.378 0.211 17.79 0.622 0.303
T-NeRF [15] 16.64 0.719 0.559 16.36 0.637 0.434 15.29 0.278 0.447 17.17 0.602 0.397
T-NeRF+B+D [15] 17.32 0.720 0.514 17.60 0.670 0.375 17.56 0.351 0.295 18.34 0.619 0.345
T-NeRF+B+D+S [15] 17.43 0.728 0.508 17.52 0.669 0.346 17.55 0.367 0.258 17.71 0.591 0.377
RoDynRF [27] 18.74 0.723 0.552 18.05 0.634 0.513 16.71 0.321 0.482 18.57 0.594 0.413
4D-GS [55] 15.35 0.686 0.432 13.40 0.544 0.506 14.48 0.201 0.336 14.37 0.513 0.366
Ours 16.45 0.749 0.487 15.32 0.626 0.363 18.03 0.434 0.209 17.17 0.624 0.288
[…]
Spin Teddy Wheel Average
Method mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$
HyperNeRF+B+D+S [35] 19.20 0.561 0.325 13.97 0.568 0.350 13.99 0.455 0.310 16.81 0.569 0.332
Nerfies+B+D+S [34] 18.38 0.585 0.309 13.65 0.557 0.372 13.82 0.458 0.310 16.45 0.570 0.339
T-NeRF [15] 16.36 0.504 0.530 13.38 0.550 0.475 14.73 0.477 0.362 15.70 0.538 0.458
T-NeRF+B+D [15] 18.95 0.541 0.468 13.99 0.573 0.435 15.69 0.541 0.301 17.06 0.573 0.390
T-NeRF+B+D+S [15] 19.16 0.567 0.443 13.71 0.570 0.429 15.65 0.548 0.292 16.96 0.577 0.379
RoDynRF [27] 17.41 0.484 0.570 14.34 0.537 0.613 15.20 0.449 0.478 17.00 0.535 0.517
4D-GS [55] 15.36 0.418 0.310 12.36 0.506 0.474 11.78 0.343 0.427 13.87 0.459 0.407
Ours 17.07 0.567 0.322 12.63 0.541 0.475 13.85 0.555 0.330 15.79 0.585 0.354
[…]
Table 8: Foreground analysis on the short NVIDIA dataset. We show additional quantitative analysis on the short NVIDIA sequences. We use masked metrics [15] with the provided foreground masks to evaluate image quality on only the dynamic part of the scenes. The three best results are highlighted in green in descending order of saturation.
Balloon1 Balloon2 Jumping Playground
Method mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$
T-NeRF [37] 17.90 0.897 0.245 19.89 0.953 0.133 15.24 0.914 0.285 14.81 0.957 0.267
D-NeRF [37] 18.75 0.907 0.294 18.23 0.950 0.203 19.23 0.941 0.160 16.65 0.970 0.227
HyperNeRF [35] 14.43 0.865 0.471 15.95 0.931 0.251 14.82 0.913 0.278 12.32 0.950 0.390
NR-NeRF [49] 15.87 0.880 0.383 20.13 0.962 0.177 16.76 0.934 0.228 13.91 0.956 0.311
NSFF [25] 19.34 0.925 0.225 22.63 0.969 0.155 20.42 0.955 0.152 20.22 0.983 0.162
DynamicNeRF [14] 19.73 0.931 0.134 23.79 0.981 0.042 20.68 0.958 0.100 21.05 0.984 0.095
RoDynRF [27] 19.70 0.928 0.147 23.05 0.977 0.046 21.54 0.962 0.093 20.29 0.984 0.097
TiNeuVox [11] 16.80 0.888 0.330 19.59 0.955 0.147 17.34 0.932 0.196 13.02 0.950 0.363
4D-GS [55] 18.72 0.911 0.247 22.21 0.973 0.066 18.60 0.938 0.173 17.23 0.976 0.156
Ours 19.76 0.930 0.159 22.14 0.973 0.056 21.07 0.960 0.080 19.76 0.982 0.107
Skating Truck Umbrella Average
Method mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$ mPSNR $\uparrow$ mSSIM $\uparrow$ mLPIPS $\downarrow$
T-NeRF [37] 22.56 0.993 0.086 18.40 0.913 0.409 16.32 0.945 0.496 17.87 0.939 0.274
D-NeRF [37] 22.05 0.993 0.090 19.84 0.923 0.312 17.34 0.953 0.365 18.87 0.948 0.236
HyperNeRF [35] 23.86 0.993 0.074 19.40 0.918 0.281 14.76 0.940 0.489 16.51 0.930 0.319
NR-NeRF [49] 27.17 0.995 0.052 18.25 0.912 0.534 16.23 0.948 0.404 18.33 0.941 0.298
NSFF [25] 31.37 0.998 0.019 26.51 0.971 0.135 19.28 0.967 0.238 22.82 0.967 0.155
DynamicNeRF [14] 37.22 0.999 0.010 26.75 0.973 0.067 19.66 0.969 0.152 24.13 0.971 0.086
RoDynRF [27] 34.37 0.997 0.035 27.29 0.978 0.061 20.85 0.970 0.134 23.87 0.971 0.087
TiNeuVox [11] 27.18 0.996 0.047 21.39 0.934 0.204 16.09 0.949 0.358 18.77 0.943 0.235
4D-GS [55] 30.99 0.997 0.048 22.77 0.947 0.207 18.36 0.956 0.259 21.27 0.957 0.165
Ours 36.38 0.999 0.014 27.16 0.975 0.063 21.71 0.975 0.087 24.00 0.971 0.081
As most natural scenes are dominated by static background, we individually assess the quality of dynamic foreground reconstruction, which is the most complex part of dynamic view synthesis from monocular inputs. To this end, we recalculate metrics on the NVIDIA sequences using masked metrics [15], based on the motion masks contained in the dataset. As shown in Tab. 8, our method still performs favorably regarding perceptual metrics, indicating that improvements were not exclusively located in the static background.
As our method can handle significantly longer and temporally denser sequences than the commonly used train/test splits of the NVIDIA dynamic scenes dataset, we further compare against DynIBaR [26], using their evaluation protocol containing $\sim 150$ training views per scene. For fairness, we also use the consistent video depth made available by the authors. Results are shown in  Tab. 9.
Table 9: Comparison on the long NVIDIA dataset. We compare to the state-of-the-art approach DynIBaR [26] on the long sequences from the NVIDIA dynamic scenes dataset. Masked metrics are calculated with respect to the dynamic foreground masks.
Method Balloon1 Balloon2 Jumping Playground Skating Truck Umbrella DynamicFace Average
PSNR $\uparrow$
DynIBaR [26] 28.99 28.90 23.74 28.32 31.50 33.75 28.25 29.18 29.08
Ours 27.32 30.09 22.67 25.51 33.29 32.14 24.76 28.18 28.00
SSIM $\uparrow$
DynIBaR [26] 0.920 0.942 0.866 0.953 0.969 0.966 0.898 0.971 0.936
Ours 0.902 0.934 0.841 0.895 0.969 0.952 0.737 0.962 0.899
LPIPS $\downarrow$
DynIBaR [26] 0.064 0.070 0.126 0.049 0.058 0.049 0.085 0.052 0.069
Ours 0.065 0.052 0.128 0.072 0.044 0.045 0.142 0.048 0.074
mPSNR $\uparrow$
DynIBaR [26] 23.91 25.53 17.49 20.91 19.52 29.06 25.96 30.55 24.12
Ours 22.35 23.51 16.30 18.30 19.89 25.57 21.75 23.94 21.45
mSSIM $\uparrow$
DynIBaR [26] 0.944 0.985 0.911 0.989 0.990 0.988 0.971 0.997 0.972
Ours 0.937 0.980 0.896 0.984 0.992 0.979 0.921 0.991 0.960
mLPIPS $\downarrow$
DynIBaR [26] 0.079 0.032 0.168 0.062 0.070 0.038 0.043 0.010 0.062
Ours 0.077 0.037 0.193 0.097 0.069 0.053 0.110 0.026 0.083
While our method does not quite attain the image quality of DynIBaR, we achieve competitive results with significantly lower training and rendering times.
Table 10: Timings for individual method components.
      Component       Time [s]
      Data Preprocessing       140
      Depth Alignment       6
      Model Initialization       1
      Model Optimization       817
We further provide a breakdown of timings for individual components of our method on the NVIDIA dataset in Tab. 10. Note that the cost of data preprocessing highly depends on the amount of input images, while the optimization and rendering times increase with image resolution.
A.4 Concurrent Work
Concurrent to our work, other methods on arXiv address dynamic reconstruction and novel-view synthesis from monocular video. Lee et al. [23] propose a new scene representation for fast reconstruction and interactive rendering from in-the-wild monocular videos. Similar to our approach, they separately model the foreground and background, but use plane based proxy geometry for static regions, and point clouds for dynamic foreground, enabling fast forward rendering and an impressive image quality. More details and results are presented at the official project page: https://casual-fvs.github.io.
Following recent trends in static scene reconstruction, Wang et al. [50] integrate diffusion-based priors in the monocular reconstruction. While not being exceptionally fast, their approach hallucinates additional training supervision for occluded areas, improving results on the challenging iPhone dataset [15]. Again, we refer the interested reader to the official project page for more information: https://mightychaos.github.io/dpdy_proj/.
A.5 Societal Impact
Our method reconstructs a 4D representation and generates novel views based on a monocular input video. As the method itself cannot fill unseen areas or use/provide any information that goes beyond the inputs provided by the users themselves, we do not expect any negative societal impact resulting from our work. In terms of environmental impact, the short training times and consumer-grade GPU (RTX 3090/4090) required by our method makes it very economic in terms of power consumption and hardware requirements: Overall, after downloading and setting up our code, running a single reconstruction, including data preprocessing with user inputs, takes under one hour of time. When running on a slightly above average PSU with $750$W at full capacity (which is usually not the case), given an average energy price of \$$0.173$ per kWh (April 2024 in the U.S.), we get a worst-case cost of \$$0.13$ per scene reconstruction. Thus, in comparison to DynIBaR [26], the current the state-of-the-art in terms of image quality, which requires $384$ GPU hours on high-end A100 GPUs, our method has a significantly lower environmental footprint.

### v2

@@ 4.2 Qualitative Results
We show qualitative comparisons on the NVIDIA (Fig. 2) and iPhone (Fig. 3) datasets. For the iPhone sequences, we show the best version of HyperNeRF [PSH21] and T-NeRF [PCPMMN21] according to image quality metrics, using all additional regularizers (Lidar depth, random background, and distortion loss) proposed by Gao et al. [GLT22]. Again, we find that our method achieves a competitive and oftentimes superior perceived image quality, preserving higher frequencies and clean edges. As our 4D model is temporally conditioned and does not directly estimate inter-frame motion, it can well handle scene movements of varying complexity (e.g., the rotating motion on the Paper-windmill scene in Fig. 3 or human motion such as in the Skating scene in LABEL:{fig:nvidia}), as long as it is constantly observed from the input camera trajectory (and not occluded).
Figure 4: Novel views synthesized using complex in-the-wild sequences of the DAVIS dataset. On the right, we show the corresponding visualizations of our implicit foreground/background separation.
Figure 4 further provides novel views synthesized using complex in-the-wild sequences from the DAVIS dataset. We additionally include visualizations of our implicit foreground/background separation obtained by only rendering the plain static or dynamic components of the sampled point clouds. Given that our method is efficient in both memory and compute, we train our DAVIS models on the full original resolution (FHD) and double the sampled point cloud size to $8M$ points to maintain a dense distribution in the rasterized feature images. This upgrade in rendering resolution and point queries increases the training time to $0.52$ hours and reduces online framerates to $18$ fps. Overall, our experiments show that DNPC robustly handles high-resolution in-the-wild inputs provided that decent camera poses can be obtained from current SfM methods. For more video results, including foreground/background separation, 3D anaglyphs renderings and live GUI capture, please see our supplemental material.
@@ 4.3 Ablation Study
Table 3: Comparison of ablated D-NPC versions in terms of image quality, model size, and speed.
@@ 5 Discussion and Limitations
The conducted experiments demonstrate the adaptability and effectiveness of the implicit neural point cloud representation [HFK24] for non-rigid novel view synthesis; Our temporal extension D-NPC effectively leverages data-driven priors for dynamic reconstruction from single monocular video, yielding a favorable trade-off between optimization time and image quality. Due to the fast forward-rendering approach, it is currently one of the only purely monocular methods enabling real-time frame rates for interactive scene exploration. Especially in terms of perceptual LPIPS and high image frequencies, our method quantitatively and visually outperforms most recent monocular view synthesis methods while only requiring a fraction of the optimization time. However, there is room for improvement: Our method relies on a decent quality of camera poses and monocular depth. Moreover, our initialization process assumes that the dynamic content is fully contained in the camera frustum, resulting in visible crops when synthesizing far-off novel views, as shown in Fig. 5. Finally, our method handles videos of up to $\approx 500$ frames (similar to related approaches) using less than $24$GB of GPU-VRAM at FHD resolution, with longer video sequences remaining a limitation.
@@ 6 Conclusion
We presented Dynamic Neural Point Clouds, an efficient implicit point cloud-driven [HFK24] approach for non-rigid novel-view synthesis from single monocular video inputs. In contrast to existing ray marching-based NeRF methods, our rasterization-based forward rendering approach enables a direct and elegant way to initialize and guide the reconstruction from data-driven priors, not only making it one of the fastest approaches in terms of optimization time, but also enabling rendering at real-time frame rates. While our experiments show our novel approach already yields competitive image quality according to several metrics on common benchmark datasets, we believe that our work constitutes a promising new direction for future research and extensions, like the integration of diffusion-based priors.
Acknowledgments
We thank Jann-Ole Henningson for help with figures and writing. This work was partially funded by the DFG under projects “Real-Action VR” (ID 523421583), and “Increasing Realism of Omnidirectional Videos in Virtual Reality” (ID 491805996); the L3S Research Center, Hanover, Germany; as well as by the ERC Consolidator Grant 4DRepLy (770784). Open Access funding enabled and organized by Projekt DEAL.